# Robust Weight Initialization for Tanh Neural Networks with Fixed Point Analysis

### v1

@@ Abstract
As a neural network’s depth increases, it can achieve strong generalization performance. Training, however, becomes challenging due to gradient issues. Theoretical research and various methods have been introduced to address this issues. However, research on weight initialization methods that can be effectively applied to tanh neural networks of varying sizes still needs to be completed. This paper presents a novel weight initialization method for Feedforward Neural Networks with tanh activation function. Based on an analysis of the fixed points of the function $\tanh(ax)$, our proposed method aims to determine values of $a$ that prevent the saturation of activations. A series of experiments on various classification datasets demonstrate that the proposed method is more robust to network size variations than the existing method. Furthermore, when applied to Physics-Informed Neural Networks, the method exhibits faster convergence and robustness to variations of the network size compared to Xavier initialization in problems of Partial Differential Equations.
@@ 1 Introduction
Deep learning has enabled substantial advancements in state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016). In general, the expressivity of neural networks exponentially increases with depth (Poole et al., 2016; Raghu et al., 2017), enabling strong generalization performance. This increased depth, though, can result in vanishing or exploding gradients and poor signal propagation throughout the model (Bengio et al., 1993), prompting the development of various weight initialization methods. Xavier initialization (Glorot & Bengio, 2010) ensures signals stay in the non-saturated region for sigmoid and hyperbolic tangent activations, while He initialization (He et al., 2015) maintains stable variance for ReLU networks. Especially in ReLU neural networks, several weight initialization methods have been proposed to mitigate the dying ReLU problem, which hinders signal propagation in deep networks (Lu et al., 2019; Lee et al., 2024). However, to the best of our knowledge, research on the initialization method to tackle the stability of extremely deep tanh networks during training is still limited. Such networks commonly use Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are widely applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
@@
The main contribution of this paper is the proposal of a simple weight initialization method for FeedForward Neural Networks (FFNNs) with tanh activation function. This method facilitates effective learning across a range of network sizes, outperforming Xavier initialization by reducing the need for extensive hyperparameter tuning such as the number of hidden layers and units. The theoretical foundation for this approach is provided through the fixed point of the function $\tanh(ax)$. We experimentally demonstrate that the proposed method achieves higher validation accuracy and lower validation loss compared to the Xavier initialization method across various FFNN network sizes on the MNIST, Fashion MNIST, and CIFAR-10 datasets. Additionally, the proposed method demonstrates its effectiveness in training across various network configurations within PINNs. Notably, while Xavier initialization shows decreasing loss as network depth increases, it fails to maintain performance beyond a certain depth, leading to increased loss and poor training outcomes. In contrast, the proposed method continues to improve performance even at greater depths, ensuring stable training and better results.
Contributions. Our contributions can be summarised as follows:
@@
  • We show the conditions under which activation values do not vanish as we increase the depth of the neural network, using a fixed-point analysis (Section 3.1 and 3.2).
@@
  • We propose a novel weight initialization method for tanh-based neural networks that has strong robustness to variations in network size (Section 3.2 and 3.3).
@@
  • We experimentally demonstrate that the proposed method is more robust to variations in network size than Xavier initialization on image benchmark datasets and PINNs (Section 4).
@@ 2 Related works
The expressivity of neural networks typically grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017). Weight initialization is crucial for training deep networks effectively (Saxe et al., 2014; Mishkin & Matas, 2016). Xavier (Glorot & Bengio, 2010) and He He et al. (2015) initialization are common initialization methods typically used with tanh and ReLU activation functions, respectively. Various initialization methods have been proposed to facilitate the training of deeper ReLU neural networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024). However, to the best of our knowledge, research on weight initialization for neural networks with tanh activation remains limited. Tanh neural networks have been increasingly used, particularly in physics-informed neural networks (PINNs).
@@
PINNs have shown promising results in solving forward, inverse, and multiphysics problems arising in science and engineering. (Lu et al., 2021; Karniadakis et al., 2021; Cuomo et al., 2022b; a; Yin et al., 2021; Wu et al., 2023; Hanna et al., 2022; Bararnia & Esmaeilpour, 2022; Shukla et al., 2020; Zhu et al., 2024; Hosseini et al., 2023; Mao et al., 2020). PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are typically trained by minimizing a loss defined by the sum of least-squares that incorporates the residual of PDE, boundary conditions, and initial conditions. This loss is usually minimized using gradient-based optimizers such as Adam (Kingma, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both. Universal approximation theories (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2020; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) guarantee the capability and performance of neural networks as an approximation of the analytic solution to PDE. However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning optimal hyperparameters of loss terms. To alleviate these issues, many authors have introduced enhanced versions of PINNs: (1) the self-adaptive loss balanced PINNs (lbPINNs) that automatically adjust the hyperparameters of loss terms during the training process (Xiang et al., 2022), (2) the Bayesian PINNs (B-PINNs) that are specialized to deal with forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs) that are trained with the gradient information from the numerical solution by the multigrid method and designed for solving stationary PDEs  (Peng et al., 2022), (4) Auxiliary Pinns (A-PINNs) that effectively handle integro-differential equations (Yuan et al., 2022), (5) conservative PINNs (cPINNs) and exetended PINNs (XPINNs) that adopt the domain decomposition technique (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) parrel PINNs that reduces the computational cost of cPINNs and XPINNs (Shukla et al., 2021), (7) gradient-enhanced PINNs (gPINNs) that use the gradient of the PDE loss term with respect to the network inputs  (Yu et al., 2022).
PINNs primarily employ Xavier initialization for training (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024), but our experimental results indicate that this method limits the performance of larger network sizes. Although there have been recent results on initialization methods for PINNs, most of them have relied on transfer learning (Tarbiyati & Nemati Saray, 2023). Thus, we propose a weight initialization method that does not require transfer learning and is robust to variations in network size.
@@
Figure 1: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using Xavier initialization (Left) and the proposed initialization (Right). Experiments were conducted on networks with 10,000 hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
@@ 3 Proposed Weight Initialization method
In this section, we discuss the proposed weight initialization method. Subsection 3.1 introduces the theoretical motivation behind the methodology. Subsection 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Subsection 3.1. Finally, in Subsection 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
@@ 3.1 Theoretical motivation
Experimental results in Figure 1 reveal that when Xavier initialization is employed in FFNNs with tanh activation, the distribution of activation values tends to cluster around zero in deeper layers. This vanishing of activation values can hinder the training process due to a discrepancy between the activation values and the desired output. However, theoretically preventing this phenomena is not straightforward. In this subsection, we gives a theoretical analysis based on a fixed point of $\tanh(ax)$ to bypass the phenomena. Before giving the theoretical foundations, consider the basic results for a tanh activation function. Recall that $x^{*}$ is a fixed point of a function $f$ if $x^{*}$ belongs to both the domain and the codomain of $f$, and $f(x^{*})=x^{*}$. The proofs of Lemma 1 and Lemma 2 are provided in Appendix A.
@@
For a fixed $a>0$ define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
@@
Then, there exists a fixed point $x^{\ast}$. Furthermore,
@@
    if $0<a\leq 1$, then $\phi$ has a unique fixed point $x^{\ast}=0$.
@@
    if $a>1$, then $\phi$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
Remark that the function $\phi_{a}$ can be considered as one-layer tanh FFNN.
@@
For a given initial value $x_{0}>0$ define
@@
Note that the parameter $a$ in Lemma 2 does not change across all iterations. In Propositions 3 and Corollary 4, we address cases where the value of $a$ varies with each iteration.
@@
Then, the inequality $\hat{\Phi}_{m}\leq\Phi_{m}\leq\tilde{\Phi}_{m}$ holds for all $m$. By Lemma 1, for any $x\geq 0$, we have $\lim_{m\to\infty}\hat{\Phi}_{m}=0$ and $\lim_{m\to\infty}\tilde{\Phi}_{m}=0$. Therefore, the Squeeze Theorem guarantees that $\lim_{m\to\infty}\Phi_{m}(x)=0$.
@@
$\left|\lim_{m\to\infty}\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}$
@@ 3.2 The derivation of the proposed weight initialization method
To establish the notation, consider a feedforward neural network with $L$ layers. The network processes $K$ training samples, denoted as pairs $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{K}$, where $\bm{x}_{i}\in\mathbb{R}^{N_{x}}$ is training input and $\bm{y}_{i}\in\mathbb{R}^{N_{y}}$ is its corresponding output. The iterative computation at each layer $\ell$ is defined as follows:
@@
where $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ is the weight matrix, $\mathbf{b}^{\ell}\in\mathbb{R}^{N_{\ell}}$ is the bias, and $\tanh(\cdot)$ is an element-wise activation hyperbolic tangent function.
We present a simplified analysis of signal propagation in FFNNs with the tanh activation function. For notational convenience, it is assumed that all hidden layers, as well as the input and output layers, have a dimension of $n$, i.e., $N_{\ell}=n$ for all $\ell$. Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as:
$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)},\,\text{for }i=1,\dots,n.$
For the $k+1$-th layer, $i=1,\ldots,n$, this expression can be generalized as:
$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}.$ (1)
According to Lemma 2, when $a>1$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$, the sequence $\{x_{k}\}$ defined by $x_{k+1}=\tanh(ax_{k})$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $k\to\infty$. This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$ and ensures that the activation values do not vanish as network depth increases. Furthermore, by Lemma 2, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, then $x_{i}^{L}$ approaches zero. Therefore, to ensure that ((((i)))) $a_{i}^{k}$ remains close to 1 and ((((ii)))) $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$, we design the initial weight matrix as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, 0 otherwise, and $\mathbf{Z}^{\ell}$ is a noise matrix drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ and $\alpha=0.085$. Then $a_{i}^{k+1}$ follows the distribution:
@@
According to Equation 2, the mean of $a_{i}^{k+1}$ is $1$, so choosing an appropriate $\sigma_{z}$ satisfies condition ((((i)))). For condition ((((ii)))), if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of $a_{i}^{k+1}$ increases, as indicated by Equation 2. As a result, the probability that the absolute value of $x_{i}^{k+1}$ surpasses that of $x_{i}^{k}$ is higher. However, if $\sigma_{z}$ is too small, the increase in the variance of $a_{i}^{k+1}$ becomes limited. Therefore, choosing an appropriate $\sigma_{z}$ is crucial.
@@
Figure 2: Difference between maximum and minimum activation values at each layer when propagating $3,000$ input data through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$. Network with $10,000$ hidden layers, each with 32 nodes (Left), and a network with alternating hidden layers of $64$ and $32$ nodes (Right).

### v2

@@ Abstract
As a neural network’s depth increases, it can improve generalization performance. However, training deep networks is challenging due to gradient and signal propagation issues. To address these challenges, extensive theoretical research and various methods have been introduced. Despite these advances, effective weight initialization methods for tanh neural networks remain insufficiently investigated. This paper presents a novel weight initialization method for neural networks with tanh activation function. Based on an analysis of the fixed points of the function $\tanh(ax)$, the proposed method aims to determine values of $a$ that mitigate activation saturation. A series of experiments on various classification datasets and physics-informed neural networks demonstrates that the proposed method outperforms Xavier initialization methods (with or without normalization) in terms of robustness across different network sizes, data efficiency, and convergence speed. Code is available at https://github.com/1HyunwooLee/Tanh-Init.
@@ 1 Introduction
Deep learning has significantly advanced state-of-the-art performance across various domains (LeCun et al., 2015; He et al., 2016). The expressivity of neural networks increases exponentially with depth, enhancing generalization performance (Poole et al., 2016; Raghu et al., 2017). However, deeper networks often face gradient and signal propagation issues (Bengio et al., 1993). These challenges have driven the development of effective weight initialization methods designed for various activation functions. Xavier initialization (Glorot & Bengio, 2010) prevents saturation in sigmoid and tanh activations, while He initialization (He et al., 2015) stabilizes variance for ReLU networks. Especially in ReLU neural networks, several weight initialization methods have been proposed to mitigate the dying ReLU problem, which hinders signal propagation in deep networks (Lu et al., 2019; Lee et al., 2024). However, to the best of our knowledge, research on initialization methods that are robust across different sizes of tanh networks is underexplored. Tanh networks commonly employ Xavier initialization (Raissi et al., 2019; Jagtap et al., 2022; Rathore et al., 2024) and are applied in various domains, such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) and Recurrent Neural Networks (RNNs) (Rumelhart et al., 1986), with performance often dependent on model size and initialization randomness (Liu et al., 2022).
@@
The main contribution of this paper is the proposal of a simple weight initialization method for FeedForward Neural Networks (FFNNs) with tanh activation function. The proposed method is data-efficient and demonstrates robustness across different network sizes. Moreover, it reduces dependency on normalization techniques such as Batch Normalization (BN) (Ioffe, 2015) and Layer Normalization (LN) (Ba, 2016). As a result, it alleviates the need for extensive hyperparameter tuning, such as selecting the number of hidden layers and units, while also eliminating the computational overhead associated with normalization. The theoretical foundation of this approach is based on the fixed point of the function $\tanh(ax)$. We evaluate the proposed method on two tasks: classification and PINNs. For classification tasks, we assess its performance across various FFNN sizes using standard benchmark datasets. The results demonstrate improved validation accuracy and lower loss compared to Xavier initialization with BN or LN. For PINNs, the method exhibits robustness across diverse network sizes and demonstrates its effectiveness in solving a wide range of PDE problems. Notably, for both tasks, the proposed method outperforms Xavier initialization in data efficiency; that is, it achieves improved performance even with limited data. Our main contributions can be summarized as follows:
@@
  • We identify the conditions under which activation values do not vanish as the neural network depth increases, using a fixed-point analysis (Section 3.1 and 3.2).
@@
  • We propose a novel weight initialization method for tanh neural networks that is robust across different network sizes and demonstrates high data efficiency (Section 3.2 and 3.3).
@@
  • We experimentally show that the proposed method is more robust across different network sizes on image benchmarks and PINNs (Section 4).
  • We experimentally show that the proposed method is more data-efficient than Xavier initialization, with or without normalization, on image benchmarks and PINNs (Section 4).
@@ 2 Related works
The expressivity of neural networks grows exponentially with depth, resulting in improved generalization performance (Poole et al., 2016; Raghu et al., 2017). Proper weight initialization is crucial for effectively training deep networks (Saxe et al., 2014; Mishkin & Matas, 2016). Xavier (Glorot & Bengio, 2010) and He He et al. (2015) initialization are common initialization methods typically used with tanh and ReLU activation functions, respectively. Several initialization methods have been proposed to improve the training of deep ReLU networks (Lu et al., 2019; Bachlechner et al., 2021; Zhao et al., 2022; Lee et al., 2024). However, research on weight initialization for tanh-based neural networks remains relatively limited. Nevertheless, tanh networks have gained increasing attention in recent years, particularly in applications such as PINNs, where their performance is highly sensitive to initialization randomness.
@@
PINNs have shown promising results in solving forward, inverse, and multiphysics problems in science and engineering. (Mao et al., 2020; Shukla et al., 2020; Lu et al., 2021; Karniadakis et al., 2021; Yin et al., 2021; Bararnia & Esmaeilpour, 2022; Cuomo et al., 2022; Hanna et al., 2022; Hosseini et al., 2023; Wu et al., 2023; Zhu et al., 2024). PINNs approximate solutions to partial differential equations (PDEs) using neural networks and are trained by minimizing a loss function, typically formulated as a sum of least-squares terms incorporating PDE residuals, boundary conditions, and initial conditions. This loss is commonly optimized using gradient-based methods such as Adam (Kingma & Ba, 2014), L-BFGS (Liu & Nocedal, 1989), or a combination of both. Universal approximation theorems (Cybenko, 1989; Hornik et al., 1989; Hornik, 1991; Park et al., 2021; Guliyev & Ismailov, 2018b; Shen et al., 2022; Guliyev & Ismailov, 2018a; Maiorov & Pinkus, 1999; Yarotsky, 2017; Gripenberg, 2003) establish the theoretical capability of neural networks to approximate analytic PDE solutions. However, PINNs still face challenges in accuracy, stability, computational complexity, and tuning appropriate hyperparameters of loss terms.
To address these challenges, various improved variants of PINNs have been proposed: (1) Self-adaptive loss-balanced PINNs (lbPINNs), which automatically adjust the hyperparameters of loss terms during training (Xiang et al., 2022), (2) Bayesian PINNs (B-PINNs), designed to handle forward and inverse nonlinear problems with noisy data (Yang et al., 2021), (3) Rectified PINNs (RPINNs), which incorporate gradient information from numerical solutions obtained via the multigrid method and are specifically designed for solving stationary PDEs (Peng et al., 2022), (4) Auxiliary PINNs (A-PINNs), developed to effectively handle integro-differential equations (Yuan et al., 2022), (5) Conservative PINNs (cPINNs) and Extended PINNs (XPINNs), which employ domain decomposition techniques (Jagtap et al., 2020; Jagtap & Karniadakis, 2020), (6) Parallel PINNs, designed to reduce the computational cost of cPINNs and XPINNs (Shukla et al., 2021), and (7) Gradient-enhanced PINNs (gPINNs), which incorporate the gradient of the PDE loss term with respect to network inputs (Yu et al., 2022).
While these advancements address various challenges in PINNs, activation functions and their initialization strategies remain crucial for achieving optimal performance.The tanh activation function has been shown to perform well in PINNs (Raissi et al., 2019), with detailed experimental results provided in Appendix C.2. Xavier initialization is widely employed as the standard approach for tanh networks in existing studies (Jin et al., 2021; Son et al., 2023; Yao et al., 2023; Gnanasambandam et al., 2023; Song et al., 2024). However, our experimental results indicate that employing Xavier initialization leads to decreased model performance as network size increases. Moreover, performance gains from Batch Normalization or Layer Normalization remain limited, and Xavier initialization exhibits sensitivity to the amount of training data, particularly in smaller datasets. Although a recent study has proposed an initialization method for PINNs, it depends on transfer learning (Tarbiyati & Nemati Saray, 2023). To address these limitations, we propose a weight initialization method that is robust across different network sizes, achieves high data efficiency, and reduces reliance on both transfer learning and normalization techniques.
@@
Figure 1: The difference between the maximum and minimum activation values at each layer when propagating $3,000$ input samples through a $10,000$-layer tanh FFNN, using Xavier initialization (left) and the proposed initialization (right). Experiments were conducted on distinct networks with $10,000$ hidden layers, each having the same number of nodes: $16$, $32$, $64$, or $128$.
@@ 3 Proposed Weight Initialization method
In this section, we discuss the proposed weight initialization method. Section 3.1 introduces the theoretical motivation behind the method. Section 3.2 presents how to derive the initial weight matrix that satisfies the conditions outlined in Section 3.1. Finally, in Section 3.3, we suggest the optimal hyperparameter $\sigma_{z}$ in the proposed method.
@@ 3.1 Theoretical motivation
Experimental results in Figure 1 indicate that when Xavier initialization is employed in FFNNs with tanh activation, the activation values tend to vanish toward zero in deeper layers. This vanishing of activation values can hinder the training process due to a discrepancy between the activation values and the desired output. However, it is not straightforward to theoretically establish the conditions for preventing this phenomenon. In this section, we present a theoretical analysis based on a fixed point of $\tanh(ax)$ to mitigate this issue. Before giving the theoretical foundations, consider the basic results for a hyperbolic tangent activation function. Recall that if $f:\mathbb{R}\to\mathbb{R}$ is a function, then an element $x^{*}\in\mathbb{R}$ is called a fixed point of $f$ if $f(x^{*})=x^{*}$.
@@
For a fixed $a>0$, define the function $\phi_{a}:\mathbb{R}\to\mathbb{R}$ given as
@@
Then, there exists a fixed point $x^{\ast}$ of $\phi_{a}$. Furthermore,
@@
    if $0<a\leq 1$, then $\phi_{a}$ has a unique fixed point $x^{\ast}=0$.
@@
    if $a>1$, then $\phi_{a}$ has three distinct fixed points: $x^{\ast}=-\xi_{a}$, 00, $\xi_{a}$ such that $\xi_{a}>0$.
Proof.
The proof is detailed in Appendix A.1. ∎
Note that $\tanh(x)<x$ for all $x>0$. However, according to Lemma 1, for $a>1$ the behavior of $\tanh(ax)$ changes. If $x>\xi_{a}$ (resp. $x<\xi_{a}$), then $\tanh(ax)<x$ (resp. $\tanh(ax)>x$). At $x=\xi_{a}$, the equality $\tanh(ax)=x$ holds. The following lemma addresses the convergence properties of iteratively applying $\tanh(ax)$ for any $x>0$.
@@
For a given initial value $x_{0}>0$, define
@@
Proof.
The proof is detailed in Appendix A.2. ∎
Note that for $a>1$ and $x_{0}>\xi_{a}>0$, the sequence $\{x_{n}\}$ satisfies $\tanh(ax_{n})>\tanh(ax_{n+1})>\xi_{a}$ for all $n\in\mathbb{N}$. Similarly, when $0<x_{0}<\xi_{a}$, the sequence satisfies $\tanh(ax_{n})<\tanh(ax_{n+1})<\xi_{a}$ for all $n\in\mathbb{N}$. Given $a>1$ and $x_{0}<0$, the sequence converges to $-\xi_{a}$ as $n\to\infty$ due to the odd symmetry of $\tanh(ax)$. According to Lemma 2, when $a>1$, the sequence $\{x_{n}\}$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $n\to\infty$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$.
Note that the parameter $a$ in Lemma 2 does not change across all iterations. Propositions 3 and Corollary 4, consider the case where the value of $a$ varies for each iteration.
@@
The proof is detailed in Appendix A.3. ∎
@@
$\lim_{m\to\infty}\left|\Phi_{m}(x)\right|\geq\xi_{1+\epsilon}.$
@@
The proof is detailed in Appendix A.4. ∎
Based on Proposition 3 and Corollary 4, if there exists a sufficiently large $N\in\mathbb{N}$ such that $a_{n}<1$ (resp. $a_{n}>1+\epsilon$) for all $n\geq N$, then for any $x_{0}\in\mathbb{R}\setminus\{0\}$, $\Phi_{m}(x_{0})\to 0$ (resp. $|\Phi_{m}(x_{0})|\geq\xi_{1+\epsilon}$) as $m\to\infty$. This result implies that if the sequence $\{\Phi_{m}\}_{m=1}^{M}$ is finite, $a_{n}$ for $N\leq n\leq M$, where $N$ is an arbitrarily chosen index close to $M$, significantly influence the values of $\Phi_{M}(x_{0})$.
@@ 3.2 The derivation of the proposed weight initialization method
Based on the theoretical motivations discussed in the previous section, we propose a weight initialization method that satisfies the following conditions in the initial forward pass:
  • (i)
    It prevents activation values from vanishing towards zero in deep layers.
  • (ii)
    It ensures that the distribution of activation values in deep layers is approximately normal.
Notations. Consider a feedforward neural network with $L$ layers. The network processes $K$ training samples, denoted as pairs $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{K}$, where $\bm{x}_{i}\in\mathbb{R}^{N_{x}}$ is training input and $\bm{y}_{i}\in\mathbb{R}^{N_{y}}$ is its corresponding output. The iterative computation at each layer $\ell$ is defined as follows:
@@
where $\mathbf{W}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ is the weight matrix, $\mathbf{b}^{\ell}\in\mathbb{R}^{N_{\ell}}$ is the bias, and $\tanh(\cdot)$ is an element-wise activation hyperbolic tangent function. Denote $\mathbf{W}^{\ell}=[w_{ij}^{\ell}]$.
Signal Propagation Analysis. We present a simplified analysis of signal propagation in FFNNs with the tanh activation function. For notational convenience, it is assumed that all hidden layers, as well as the input and output layers, have a dimension of $n$, i.e., $N_{\ell}=n$ for all $\ell$. Given an arbitrary input vector $\bm{x}=(x_{1},\ldots,x_{n})$, the first layer activation $\bm{x}^{1}=\tanh(\mathbf{W}^{1}\bm{x})$ can be expressed component-wise as
$x^{1}_{i}=\tanh\big{(}w^{1}_{i1}x_{1}+\cdots+w^{1}_{in}x_{n}\big{)}=\tanh\bigg{(}\Big{(}w^{1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{1}_{ij}x_{j}}{x_{i}}\Big{)}x_{i}\bigg{)}\text{ for }i=1,\dots,n.$
For the ($k+1$)-th layer this expression can be generalized as
$x_{i}^{k+1}=\tanh\left(a_{i}^{k+1}x_{i}^{k}\right),\,\text{where}\quad a_{i}^{k+1}=w^{k+1}_{ii}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\frac{w^{k+1}_{ij}x_{j}^{k}}{x_{i}^{k}}\quad\text{for }i=1,\ldots,n.$ (1)
Equation (1) follows the form of $\tanh(ax)$, as discussed in Section 3.2. According to Lemma 2, when $a>1$, for an arbitrary initial value $x_{0}>0$ or $x_{0}<0$, the sequence $\{x_{k}\}$ defined by $x_{k+1}=\tanh(ax_{k})$ converges to $\xi_{a}$ or $-\xi_{a}$, respectively, as $k\to\infty$. This result indicates that the sequence converges to the fixed point $\xi_{a}$ regardless of the initial value $x_{0}$. From the perspective of signal propagation in tanh-based FFNNs, this ensures that the activation values do not vanish as the network depth increases. Furthermore, by Proposition 3, if $a^{k}_{i}\leq 1$ for all $N\leq k\leq L$, where $N$ is an arbitrarily chosen index sufficiently close to $L$, the value of $x_{i}^{L}$ approaches zero. Therefore, to satisfy condition ((((i)))), $a_{i}^{k}$ remains close to 1, and the inequality $a_{i}^{k}\leq 1$ does not hold for all $N\leq k\leq L$.
Proposed Weight Initialization. The proposed initial weight matrix is defined as $\mathbf{W}^{\ell}=\mathbf{D}^{\ell}+\mathbf{Z}^{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$, where $\mathbf{D}^{\ell}_{i,j}=1$ if $i\equiv j\pmod{N_{\ell-1}}$, and 0 otherwise (Examples of $\mathbf{D}^{\ell}$ are provided in Appendix D). The noise matrix $\mathbf{Z}^{\ell}$ is drawn from $\mathcal{N}(0,\sigma_{z}^{2})$, where $\sigma_{z}$ is set to $\alpha/\sqrt{N^{\ell-1}}$ with $\alpha=0.085$. Then $a_{i}^{k+1}$ follows the distribution:
@@
According to Equation (2), $a_{i}^{k+1}$ follows a Gaussian distribution with a mean of $1$. Additionally, if $x_{i}^{k}$ becomes small relative to other elements in $\bm{x}^{k}$, the variance of the distribution in (2) increases. Consequently, the probability that the absolute value of $x_{i}^{k+1}$ exceeds that of $x_{i}^{k}$ becomes higher. Figure 1 (b) shows that activation values maintain consistent scales in deeper layers.
@@
Figure 2: The difference between the maximum and minimum activation values at each layer when propagating $3,000$ input samples through a $10,000$-layer tanh FFNN, using the proposed initialization with $\alpha$ set to $0.04,0.085,0.15,$ and $0.5$. Network with $10,000$ hidden layers, each with 32 nodes (left), and a network with alternating hidden layers of $64$ and $32$ nodes (right).